### v1

@@ 4 Conclusion
This paper has extended the concept of outlier detection to panel data. Outlier detection is of vital importance to many panel data applications, including data quality. Since no distributional assumptions are made, loss functions are developed to detect these outliers. Thus, only the possible presence of an outlier can be determined. These loss functions have been specialized to the cases of multiple dates and nominal time. The direction of the difference can also be accounted for. Several applications have been discussed. These include showing a general approach to using loss functions, developing loss functions from preexisting outlier criteria, using loss functions with GIS, and developing a specialized loss function for variables classified by functions of their successive values.

### v2

@@ II.3 Figures
We have extended the concept of outlier detection to panel data. Outlier detection is of vital importance to many panel data applications, including data quality. Since no distributional assumptions are made, loss functions are developed to detect these outliers. Thus, only the possible presence of an outlier can be determined. Time can either be chronological or nominal. The direction of the difference can be accounted for. We developed loss functions for nonnegative data in Part I. We discussed everal applications. These include showing a general approach to using loss functions, developing loss functions from preexisting outlier criteria, using loss functions with GIS, developing a time-invariant loss function to apply to multiple dates at once and developing a specialized loss function for variables classified by functions of their successive values. In Part II, we dropped the sign restriction to allow any real data. The geometry restricts the choice of loss function parameter at the cost of allowing the data to inform its choice.
Authors’ Contributions
Bryan developed several of the ideas. Coleman performed all derivations and created all text and graphics.
Supplementary Materials
The R package Outliers in https://github.com/chuckcoleman/Outliers contains R code to create the Maps in Part I and Octave code to create the Figures in Part II.